Let P be a convex lattice polygon with one interior lattice point. The number of lattice points on the boundary of P plus the corresponding number for the dual polygon (to be defined) always equals 12. We will sketch four different proofs of this result, and explain why this 12 is the same as the 12 that arises in various other branches of mathematics.